# Middleware for LLMs: Tools Are Instrumental for Language Agents in Complex Environments

### v1

@@ Abstract
The applications of large language models (LLMs) have expanded well beyond the confines of text processing, signaling a new era where LLMs are envisioned as generalist language agents capable of operating within complex real-world environments. These environments are often highly expansive, making it impossible for the LLM to process them within its short-term memory. Motivated by recent research on extending the capabilities of LLMs with tools, this paper investigates the intriguing potential of tools to augment LLMs in handling such complexity. To this end, we design customized tools to aid in the proactive exploration within these massive environments. Such tools can serve as a middleware layer shielding the LLM from environmental complexity. In two representative complex environments—knowledge bases (KBs) and databases—we demonstrate the significant potential of augmenting language agents with tools in complex environments. Notably, equipped with these tools, GPT-4 achieves 2.8$\times$ the performance of the best baseline in tasks requiring access to database content and 2.2$\times$ in KB tasks. Our findings illuminate the path for advancing language agents in complex real-world applications.  [^1] Our code and data will be released at OSU_NLP/Fuxi.
@@ 1 Introduction
Large language models (LLMs) have demonstrated revolutionary language capabilities, demonstrating a human-like mastery over text OpenAI (2023a, b); Touvron et al. (2023); Jiang et al. (2024). However, the true ambition of AI extends well beyond the realm of text. The goal is to ultimately empower LLMs to act as generalist language agents that can aid humans across the multitude of complex real-world tasks Yao et al. (2022); Schick et al. (2023); Gu et al. (2023), which often involve handling complex environments, whether it be browsing intricate webpages Deng et al. (2023) or managing vast databases with millions of entries Li et al. (2023a).
@@
For LLMs to effectively serve as agents that ground human instructions into accurate actions within the environment, they must develop a robust understanding of the environment. The most direct method to achieve it is to linearize the environment into a sequence of tokens that fit into the LLM’s short-term memory (i.e., its input window) and have the LLM process the environment based on the linearized description Tai et al. (2023); Shridhar et al. (2021); Liu et al. (2023). However, such a method faces steep challenges in scaling to more complex environments, primarily due to the input size limitations of LLMs. Also, discrete token descriptions may not reflect the most natural perception of the environment. Recent work has explored using tools to extend the boundary of the LLM’s capacity Li et al. (2023b); Qin et al. (2023b); Schick et al. (2023). The core idea is that LLMs can actively decide a proper tool to use, using language as a powerful vehicle of thought Su (2023). For example, the LLM may invoke a calculator when facing a computationally intensive math task. Intuitively, we can also equip the LLM with tools that enable navigating complex environments, so that the LLM can proactively invoke different tools to explore the environment, thus circumventing limitations posed by its short-term memory (Figure 1). However, this promising paradigm has been thus far underexplored. In this paper, we aim to delve into this paradigm and answer an intriguing question: How effectively can LLMs handle complex environments with the aid of tools?
@@
Answering this question requires equipping the LLM with a suite of tools designed to meet a wide range of needs within the target environment. In this paper, we carefully develop such tailored tools for two exemplar complex environments, i.e., databases and knowledge bases (KBs). Unlike readily available Web APIs Qin et al. (2023b) used in prior research, our tools have to be manually invented from scratch. In crafting these tools, we capitalize on the intuition of human information-gathering behaviors—such as performing keyword searches to identify a relevant database column or investigating the connections of a KB entity—to fulfill complex tasks in these intricate environments (Section 3.1). Ideally, these tools are designed to function as a middleware layer between the LLM and the environment, shielding the LLM from environmental complexity. With these specialized tools in place, we adapt ReAct Yao et al. (2022), a standard framework that enables the LLM to synergistically combine reasoning with tool usage, as our reasoning algorithm to allow the LLM to effectively leverage the provided tools (Section  3.2). The combination of the crafted tools and the reasoning algorithm allows the LLM to actively explore the environment and ground human instructions into accurate actions. We call this framework Fuxi (i.e., flexible grounding with exploration).
@@
With Fuxi, we evaluate different LLMs on benchmarks featuring complex tasks over the target environments, including a newly curated benchmark for the KB. The outcomes of our experiments are revealing: LLMs equipped with customized tools demonstrate a significant enhancement in their ability to engage with complex environments, markedly surpassing the prior art. In particular, despite its simplicity, Fuxi allows GPT-4 OpenAI (2023a) to achieve 2.8$\times$ the performance (i.e., $38.3$% vs. $13.8$%) of the best baseline in tasks requiring access to database content and 2.2$\times$ (i.e., $59.3$% vs. $27.1$%) in KB tasks. Our findings underscore the integral role of tool augmentation in enabling LLMs to handle complex environments.
@@
Our main contributions are as follows: a) We develop Fuxi, a new framework with customized tools for two complex environments, to investigate the role of tools in handling complex environments with LLMs; b) We extensively evaluate six different LLMs on our carefully chosen benchmarks; c) Our analysis highlights a critical takeaway: augmenting LLMs with tools is crucial for successfully tackling complex environments, opening new possibilities to progress LLMs as generalist language agents for practical applications.
@@ 2 Related Work
    Existing methods that feed the environment directly into the LLM for grounding Chandu et al. (2021) would fail in complex environments due to scalability issues. Specifically, these methods process the environment by linearizing it into discrete tokens Hwang et al. (2019); Shridhar et al. (2021); Yu et al. (2023); Liu et al. (2023); Tai et al. (2023); Song et al. (2023). However, linearizing expansive environments like databases with millions of entries Li et al. (2023a) or lengthy webpage HTML code Deng et al. (2023) can often exceed an LLM’s input length constraints. Alternative studies bypass the LLM’s direct interaction with complex environments by generating ungrounded draft plans for post-processing grounding Li et al. (2023c); Nie et al. (2023) or by using the LLM to assess grounded plans created via predefined rules Gu et al. (2023). Such strategies do not fully utilize the LLMs’ innate reasoning potential in actively navigating complex environments. In this paper, we explore a new paradigm where we can bypass these issues by equipping LLMs with a suite of comprehensive tools to actively gather necessary information about the environment upon demand, leveraging the LLMs’ inherent reasoning capabilities. The most closely related work to ours is StructGPT Jiang et al. (2023b). However, the narrow tool selection of StructGPT (i.e., only two tools for KBs and three schema-level tools for databases) largely constrains its flexibility in perceiving the complex environment when handling diverse tasks.
@@
    Tools are essential for enhancing the capabilities of LLMs Schick et al. (2023); Qin et al. (2023a); Mialon et al. (2023); Hao et al. (2023). Existing research, such as ToolLLM Qin et al. (2023b) and API-Bank Li et al. (2023b), focuses on open-domain applications with a wide array of readily available RESTful APIs. In contrast, this paper specifically aims to study the potential of tools in augmenting LLMs to effectively execute tasks within complex environments, where we carefully craft the specialized tools for different environments by ourselves. In addition, research focusing on RESTful APIs typically displays shallow reasoning, while practical tasks within a complex environment typically entail a long sequence of actions (e.g., querying a KB or browsing a webpage). To enable tool use in more intricate settings within a more specific complex environment, StructGPT Jiang et al. (2023b) employs a predefined sequence of tool invocations; Chameleon Lu et al. (2023) functions in an open-loop setting where the LLM directly produces a sequence for tool usage before any execution occurs. Both of them fail to seamlessly integrate the reasoning capacity of the LLM with the use of tools. In this paper, we build on ReAct to tightly synergize the generation of a reasoning step and corresponding tool use. Additionally, we introduce two simple yet effective strategies aimed at improving the accuracy of action prediction.
3 Fuxi
Fuxi equips LLMs with a suite of tools specifically tailored to support an extensive variety of operations and cater to the diverse needs within a complex environment $\mathcal{E}$. These tools can serve as a feature-rich middleware layer between the LLM and $\mathcal{E}$, abstracting the LLM from having to directly interact with all of its intricacies (Section 3.1). Furthermore, to fully unleash the inherent planning capabilities of LLMs in invoking proper tools, Fuxi builds on ReAct Yao et al. (2022) to seamlessly integrate chain-of-thought (CoT) reasoning Wei et al. (2022) with tool use, with novel strategies to enhance action accuracy (Section 3.2). This unified framework allows us to reliably investigate the potential of LLMs in handling complex environments with the aid of tools.
@@ 3.2 Reasoning with Tools
We leverage ReAct to enable the LLM to effectively invoke our crafted tools. Unlike existing methods relying on rigid, human-defined workflows that follow fixed tool usage sequences  Jiang et al. (2023b), ReAct allows the LLM autonomy in proactively determining tool selection using CoT. Thus, ReAct allows us to tap into the full potential of the LLM’s reasoning capabilities.
@@
$\hat{a}_{t}$ is the concatenation of a rationale $r_{t}$ (i.e., a thought in CoT) and a concrete tool use $a_{t}$ (e.g., in Figure 2, $\hat{a}_{1}$ is the concatenation of Thought 1 and Act 1), while $o_{t}$ is an observation from the environment (i.e., the execution result of $a_{t}$). In ReAct, the LLM jointly decodes $\hat{a}_{t}$ based on $c_{t}$ for each step. However, originally designed for simpler tools like the Wikipedia Search API, ReAct is more susceptible to producing an invalid $a_{t}$ that is unfaithful to $r_{t}$ when applied to more nuanced tool usage. We propose two simple strategies to remedy this issue. The first strategy is to simply amplify ReAct by providing detailed error feedback in case of incorrect tool usage by the LLM, followed by a prompt to retry based on these messages (see Figure 2(a)). [^2] For databases, we directly use the error message from sqlite3. For KBs, we manually define several simple templates for error feedback along with each tool. This relies on the LLM’s capacity for self-correction through feedback Gou et al. (2023); Chen et al. (2023), which may not always be reliable when the underpinning LLM is weak, potentially leading to the repetition of the same mistakes Guan et al. (2023). Additionally, we present decoupled generation, where the LLM’s policy $\pi$ is split into two sequential phases (i.e., $\pi\propto\pi_{1}\circ\pi_{2}$), allowing for more nuanced control of its actions. Initially, the LLM only decodes a thought $r_{t}$ following $\pi_{1}(r_{t}|c_{t})$. Subsequently, the LLM predicts an action $a_{t}$ in a separate context, incorporating both the thought $r_{t}$ and a set of simple rules $\mathcal{M}$ that determines permissible actions of this step. This is further guided by $\pi_{2}$, formulated as $a_{t}\sim\pi_{2}(a_{t}|r_{t},\mathcal{M})$. $\mathcal{M}$ encapsulates the governing rules of the environment (e.g., the relation argument for get_neighbors() must be derived from the output of get_relations(), which is applied to the specified entity argument in prior steps), infusing prior knowledge into the LLM’s decision-making process (see Figure 2(b)). The concrete prompts used by us are shown in Appendix C.
@@
Fuxi (error feedback)
@@
Fuxi (error feedback)
@@
Fuxi (decoupled generation)
@@ 4 Benchmarks
    We curate KBQA-Agent, a new test set sourcing from existing KBQA datasets that contain complex questions. In particular, we selected $500$ diverse questions that involve at least three relations, or two relations coupled with an aggregation function (i.e., Counting or Superlative). For each question, we annotate it with a ground truth sequence of actions based on the toolset defined by us. [^3] We leverage the gold S-expressions provided by Gu and Su (2022). Our dataset has been served as part of AgentBench Liu et al. (2023). Specifically, KBQA-Agent comprises questions from three KBQA datasets on Freebase: GrailQA Gu et al. (2021), ComplexWebQ Talmor and Berant (2018), and GraphQ Su et al. (2016), ensuring a wide range of question types and sources. KBQA-Agent is designed to be more representative of challenging, real-world scenarios compared to existing benchmarks (Appendix B). It offers an ideal testbed for evaluating language agents in interacting with massive KBs. We assess this through two metrics: F1 of answer entities and Validity (VA), a binary metric evaluating the LLM’s ability to find an answer, whether correct or not.
@@ 5.1 Setup
    For the underlying LLMs, we primarily compare Fuxi with baseline methods using two of the most advanced LLMs—GPT-3.5-turbo OpenAI (2023b) and GPT-4 OpenAI (2023a)—since our goal is investigating the full potential of tool-enhanced LLMs operating within complex environments. In addition, we also explore four open-source LLMs to more comprehensively evaluate our framework: Llama2-7B-Chat, Llama2-13B-Chat Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023a), and Mixtral 8$\times$7B-Instruct-v0.1 Jiang et al. (2024).
@@
    To fully understand the potential of tool augmentation for assisting LLMs in handling complex environments, we compare Fuxi against an array of strong baselines. For text-to-SQL parsing, LLMs demonstrate a strong ability to compose SQL queries when properly prompted with the database schema (i.e., API docs prompting Rajkumar et al. (2022)). This also represents the current state-of-the-art prompting-based method when oracle knowledge is not available on Bird’s leaderboard. [^4] https://bird-bench.github.io In addition, we experiment with StructGPT Jiang et al. (2023b), which represents an advanced text-to-SQL parsing method leveraging tools. For all methods on text-to-SQL parsing, we adopt the zero-shot setting. Unlike text-to-SQL parsing, directly prompting LLMs does not generate reasonable outputs for KBQA due to the massive size of the KB schema. Instead, existing KBQA methods based on LLMs typically follow two paradigms: either first generating an ungrounded program and then grounding the program to the KB schema afterwards Li et al. (2023c); Nie et al. (2023), or gradually constructing a complex program and grounding it step by step Gu et al. (2023). We compare Fuxi with the most representative work from each paradigm, namely KB-Binder Li et al. (2023c) and Pangu Gu et al. (2023). We also include StructGPT as an additional baseline for tool use. For all KBQA methods except KB-Binder, we provide a one-shot demo to obtain more meaningful results.

### v2

@@ Abstract
The applications of large language models (LLMs) have expanded well beyond the confines of text processing, signaling a new era where LLMs are envisioned as generalist agents capable of operating within complex environments. These environments are often highly expansive, making it impossible for the LLM to process them within its short-term memory. Motivated by recent research on extending the capabilities of LLMs with tools, we seek to investigate the intriguing potential of tools to augment LLMs in handling such complexity by introducing a novel class of tools, termed middleware, to aid in the proactive exploration within these massive environments. Such specialized tools can serve as a middleware layer shielding the LLM from environmental complexity. In two representative complex environments—knowledge bases (KBs) and databases—we demonstrate the significant potential of augmenting language agents with tools in complex environments. Notably, equipped with the middleware, GPT-4 achieves 2.8$\times$ the performance of the best baseline in tasks requiring access to database content and 2.2$\times$ in KB tasks. Our findings illuminate the path for advancing language agents in real-world applications.  [^1] GitHub repo: OSU_NLP/Middleware
      Hugging Face dataset: KBQA-Agent
@@ 1 Introduction
Large language models (LLMs) have demonstrated a human-like mastery over text OpenAI (2023a, b); Touvron et al. (2023); Jiang et al. (2024). However, the true ambition of AI extends well beyond the realm of text. The goal is to ultimately empower LLMs to act as generalist language agents that can aid humans across the multitude of complex real-world tasks Yao et al. (2022); Schick et al. (2023); Liu et al. (2023), which often involve handling complex environments, be it browsing complex webpages Deng et al. (2023), managing vast databases with millions of entries Li et al. (2023a), or querying huge KBs Gu et al. (2023).
@@
For LLMs to effectively serve as agents that ground human instructions accurately within the environment, they must develop a robust understanding of the environment. The most direct method to achieve it is to linearize the environment into a sequence of tokens that fit into the LLM’s short-term memory (i.e., its input window) and have the LLM process the environment based on the linearized description Tai et al. (2023); Shridhar et al. (2021); Liu et al. (2023). However, such a method faces steep challenges in scaling to more complex environments. Also, discrete token descriptions may not reflect the most natural perception of the environment. Recent work has explored using tools to extend the boundary of the LLM’s capacity Li et al. (2023b); Qin et al. (2023b); Schick et al. (2023). The core idea is that LLMs can actively decide a proper tool to use, using language as a powerful vehicle of thought Su (2023). Intuitively, we can also equip the LLM with tools that enable navigating complex environments, so that the LLM can proactively invoke different tools to explore the environment, thus circumventing limitations posed by its short-term memory (Figure 1). However, this promising paradigm has been thus far underexplored. In this paper, we aim to delve into this paradigm and answer an intriguing question: How effectively can LLMs handle complex environments with the aid of tools?
@@
Answering this question requires equipping the LLM with a suite of tools designed to meet a wide range of needs within the target environment. In this paper, we carefully develop such tailored tools for two exemplar complex environments, i.e., databases and knowledge bases (KBs). Unlike readily available Web APIs Qin et al. (2023b) used in prior research, our tools have to be manually invented from scratch. In crafting these tools, we capitalize on the intuition of human information-gathering behaviors—such as performing keyword searches to identify a relevant database column or investigating the connections of a KB entity—to fulfill complex tasks in these environments (Section 3.1). Ideally, these tools are designed to function as a middleware layer between the LLM and the environment, shielding the LLM from environmental complexity. With these specialized tools, we propose two novel schemes to enable the LLM to more accurately orchestrate its internal reasoning and tool usage: error feedback and decoupled generation (Section 3.2). The combination of the crafted tools and the tool-use schemes allows the LLM to actively explore the environment and ground human instructions into accurate actions.
@@
We evaluate different LLMs on benchmarks featuring complex tasks over the target environments, including a newly curated benchmark for the KB. The results are revealing: LLMs equipped with customized tools demonstrate a significant enhancement in their ability to engage with complex environments, markedly surpassing the prior art. In particular, despite its simplicity, such a middleware layer allows GPT-4 OpenAI (2023a) to achieve 2.8$\times$ the performance (i.e., $38.3$% vs. $13.8$%) of the best baseline in tasks requiring access to database content and 2.2$\times$ (i.e., $59.3$% vs. $27.1$%) in KB tasks. Our findings underscore the integral role of tool augmentation in enabling LLMs to handle complex environments.
@@
Our main contributions are as follows: a) We develop a new framework with customized tools for two complex environments, to investigate the role of tools in handling complex environments with LLMs; b) We evaluate six different LLMs on our carefully chosen benchmarks for a comprehensive analysis; c) Our analysis highlights a critical takeaway: augmenting LLMs with tools is crucial for successfully tackling complex environments, opening new possibilities to progress LLMs as generalist language agents for practical applications.
@@ 2 Related Work
    Existing methods that feed the environment directly into the LLM for grounding Chandu et al. (2021) would fail in complex environments due to scalability issues. Specifically, these methods process the environment by linearizing it into discrete tokens Hwang et al. (2019); Shridhar et al. (2021); Yu et al. (2023); Liu et al. (2023); Tai et al. (2023); Song et al. (2023). However, linearizing expansive environments like databases with millions of entries Li et al. (2023a) or lengthy webpage HTML code Deng et al. (2023) can often exceed an LLM’s input length constraints. Alternative studies bypass the LLM’s direct interaction with complex environments by generating ungrounded draft plans for post-processing grounding Li et al. (2023c); Nie et al. (2023) or by using the LLM to assess grounded plans created via predefined rules Gu et al. (2023). Such strategies do not fully utilize the LLMs’ innate reasoning potential in actively navigating complex environments. In this paper, we explore a new paradigm where we can bypass these issues by equipping LLMs with a suite of comprehensive tools to actively gather necessary information about the environment upon demand, leveraging the LLMs’ inherent reasoning capabilities.
@@
    Tools are essential for enhancing the capabilities of LLMs Schick et al. (2023); Qin et al. (2023a); Mialon et al. (2023); Hao et al. (2023). Existing research, such as ToolLLM Qin et al. (2023b) and API-Bank Li et al. (2023b), focuses on open-domain applications with a wide array of readily available RESTful APIs. In contrast, this paper specifically aims to study the potential of tools in augmenting LLMs to effectively execute tasks within complex environments, where we carefully craft the specialized tools for different environments by ourselves. In addition, research focusing on RESTful APIs typically displays shallow reasoning, while practical tasks within a complex environment typically entail a long sequence of actions (e.g., querying a KB or browsing a webpage). To enable tool use in more intricate settings within a more specific complex environment, StructGPT Jiang et al. (2023b) employs a predefined sequence of tool invocations; Chameleon Lu et al. (2023) functions in an open-loop setting where the LLM directly produces a sequence for tool usage before any execution occurs. Both of them fail to seamlessly integrate the reasoning capacity of the LLM with the use of tools. In this paper, we propose two novel schemes—error feedback and decoupled generation to more seamlessly and accurately orchestrate the LLM’s internal reasoning and tool usage.
3 Middleware for LLMs
We equip LLMs with a suite of tools specifically tailored to support an extensive variety of operations and cater to the diverse needs within a complex environment $\mathcal{E}$. We call these tools middleware, as they can serve as a feature-rich middle layer between the LLM and $\mathcal{E}$, abstracting the LLM from having to directly interact with all of its intricacies (Section 3.1). Furthermore, to fully unleash the inherent reasoning capabilities of LLMs in invoking proper tools, we propose two novel schemes to enhance tool use accuracy: error feedback, which provides concrete tool use error information and expects the LLM to correct the error autonomously, and decoupled generation, where the LLM’s reasoning steps and tool use are separated for better controllability (Section 3.2). This unified framework allows us to reliably investigate the potential of LLMs in handling complex environments with the aid of tools.
@@ 3.2 Reasoning with Tools
We first choose ReAct Yao et al. (2022) to serve as the backbone of our reasoning framework, in which the LLM can proactively decide which tool to use based on its own chain-of-thought Wei et al. (2022). Based on this backbone, we propose two novel schemes to improve the accuracy of tool use, which is critical for complex tasks where successful tool use necessitates careful tool selection and precise argument assignment. Unlike existing methods relying on human-defined workflows that follow fixed-order tool usage Jiang et al. (2023b), our framework allows the LLM autonomy in proactively determining tool selection using CoT.
@@
$\hat{a}_{t}$ is the concatenation of a rationale $r_{t}$ (i.e., a thought in CoT) and a concrete tool use $a_{t}$ (e.g., in Figure 2, $\hat{a}_{1}$ is the concatenation of Thought 1 and Act 1), while $o_{t}$ is an observation from the environment (i.e., the execution result of $a_{t}$). In ReAct, the LLM jointly decodes $\hat{a}_{t}$ based on $c_{t}$ for each step. However, originally designed for simpler tools like the Wikipedia Search API, the naive ReAct framework is more susceptible to producing an invalid $a_{t}$ that is unfaithful to $r_{t}$ when applied to more nuanced tool usage. We propose two simple strategies to remedy this issue. The first strategy is to simply amplify ReAct by providing detailed error feedback in case of incorrect tool usage by the LLM, followed by a prompt to retry based on these messages (see Figure 2(a)). [^2] For databases, we directly use the error message from sqlite3. For KBs, we manually define several simple templates for error feedback along with each tool. This relies on the LLM’s capacity for self-correction through feedback Gou et al. (2023); Chen et al. (2023), which may not always be reliable when the underpinning LLM is weak, potentially leading to the repetition of the same mistakes Guan et al. (2023). Additionally, we present decoupled generation, where the LLM’s policy $\pi$ is split into two sequential phases (i.e., $\pi\propto\pi_{1}\circ\pi_{2}$), allowing for more nuanced control of its actions. Initially, the LLM only decodes a thought $r_{t}$ following $\pi_{1}(r_{t}|c_{t})$. Subsequently, the LLM predicts an action $a_{t}$ in a separate context, incorporating both the thought $r_{t}$ and a set of simple rules $\mathcal{M}$ that determines permissible actions of this step. This is further guided by $\pi_{2}$, formulated as $a_{t}\sim\pi_{2}(a_{t}|r_{t},\mathcal{M})$. $\mathcal{M}$ encapsulates the governing rules of the environment (e.g., the relation argument for get_neighbors() must be derived from the output of get_relations(), which is applied to the specified entity argument in prior steps), infusing prior knowledge into the LLM’s decision-making process (see Figure 2(b)). The concrete prompts used by us are shown in Appendix C.
@@
Middleware (error feedback)
@@
Middleware (error feedback)
@@
Middleware (decoupled generation)
@@ 4 Benchmarks
    We curate KBQA-Agent, a new test set sourcing from existing KBQA datasets that contain complex questions. In particular, we selected $500$ diverse questions that involve at least three relations, or two relations coupled with an aggregation function (i.e., Counting or Superlative). For each question, we annotate it with a ground truth sequence of actions based on the toolset defined by us. [^3] We leverage the gold S-expressions provided by Gu and Su (2022). Specifically, KBQA-Agent comprises questions from three KBQA datasets on Freebase: GrailQA Gu et al. (2021), ComplexWebQ Talmor and Berant (2018), and GraphQ Su et al. (2016), ensuring a wide range of question types and sources. KBQA-Agent is designed to be more representative of challenging, real-world scenarios compared to existing benchmarks (Appendix B). It offers an ideal testbed for evaluating language agents in interacting with massive KBs. We assess this through two metrics: F1 of answer entities and Validity (VA), a binary metric evaluating the LLM’s ability to find an answer, whether correct or not.
@@ 5.1 Setup
    For the underlying LLMs, we primarily compare Middleware with baseline methods using two of the most advanced LLMs—GPT-3.5-turbo-0613 OpenAI (2023b) and GPT-4-0613 OpenAI (2023a)—since our goal is investigating the full potential of tool-enhanced LLMs operating within complex environments. In addition, we also explore four open-source LLMs to more comprehensively evaluate our framework: Llama2-7B-Chat, Llama2-13B-Chat Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023a), and Mixtral 8$\times$7B-Instruct-v0.1 Jiang et al. (2024).
@@
    To fully understand the potential of tool augmentation for assisting LLMs in handling complex environments, we compare Middleware against an array of strong baselines. For text-to-SQL parsing, LLMs demonstrate a strong ability to compose SQL queries when properly prompted with the database schema (i.e., API docs prompting Rajkumar et al. (2022)). This also represents the current state-of-the-art prompting-based method when oracle knowledge is not available on Bird’s leaderboard. In adddition, we also compare with more baselines on Bird’s leaderboard that originally did not submit their results using no oracle knowledge (See Appendix D.1). For all methods on text-to-SQL parsing, we adopt the zero-shot setting. Unlike text-to-SQL parsing, directly prompting LLMs does not generate reasonable outputs for KBQA due to the massive size of the KB schema. Instead, existing KBQA methods based on LLMs typically follow two paradigms: either first generating an ungrounded program and then grounding the program to the KB schema afterwards Li et al. (2023c); Nie et al. (2023), or gradually constructing a complex program and grounding it step by step Gu et al. (2023). We compare Middleware with the most representative work from each paradigm, namely KB-Binder Li et al. (2023c) and Pangu Gu et al. (2023). We also include StructGPT as an additional baseline for tool use. For all KBQA methods except KB-Binder, we provide a one-shot demo to obtain more meaningful results.